20 ml of K4Fe(CN)6  solution is divided into 2 equal parts.One part is completely reacted with 10 ml FeCl3 solution to form KFe [Fe (CN)6] and KCl.Another part of K4Fe(CN) 6 solution is completely reacted with 20 ml of 0.5M FeCl2 solution to produce K2Fe [Fe(CN)6] and KCl.Molarity of FeCl3 solution is?

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
we have
K4Fe(CN)6 +FeCl3 ----KFe[Fe(CN)6] +3KCl -------1
K4Fe(CN)6 +FeCl2 ----K2Fe[Fe(CN)6] +2KCl -------2
let the molarity of FeCl3 be M
and the molarity of K4[Fe(CN)6] be M1
now from 2
mole of K4[Fe(CN)6]=mole of FeCl2
M1*10=20*.5
M1=1Molar
now apllying mole equivalence to 1
M*10=1*10
M=1 Molar
Thus the molarity of FeCl3 is 1 M